OpenStudy (anonymous):

find the ratio in simplest form: 15 minutes to 2 hours

OpenStudy (anonymous):

1 hour is 60 minutes, so 2 hours is 120 minutes, can you take it from there?

OpenStudy (anonymous):

\[\frac{(15/60)}{2}=\frac{1}{8} \]15 Minutes had to be changed to hours.

OpenStudy (anonymous):

1/8 is what i got as well... i wanted to see if he/she could figure it out.

OpenStudy (anonymous):

You should get a medal.

OpenStudy (anonymous):

Why not you? Lol.

OpenStudy (anonymous):

No problem.